# An adaptation of InfoMap to absorbing random walks using absorption-scaled graphs

### v3

@@ 6.2 Increasing node-absorption rates of bridging nodes instead of removing community bridges
6.2.1 The networks in our SIR simulations
@@
3:  Select $n_{\mathrm{WS}}\times N_{\mathrm{WS}}$ pairs $\{i_{1},i_{2}\}$ of nodes uniformly at random from all node pairs that belong to distinct ring-lattice graphs. Add bidirectional edges between the nodes of each pair. (The edges $(i_{1},i_{2})$ and $(i_{2},i_{1})$ are “community bridges”, and the nodes $i_{1}$ and $i_{2}$ are “bridging nodes”.)
@@ 6.2.2 Duration, final size, and peak of SIR outbreaks
We use Algorithm 3 to run SIR simulations on the graph $G^{\mathrm{RL}}$ for different node-absorption rates (which correspond to the recovery intensities) and transmission intensities. We refer to a set of node-absorption rates $\delta_{1},\ldots,\delta_{N}$ and transmission intensities $\beta_{1},\ldots,\beta_{N}$ of the $N$ nodes of $G^{\textnormal{RL}}$ as a parameter configuration. In steps 1 and 2 of Algorithm 3, we recursively determine parameter configurations. We refer to each step of this recursive process as a stage. For the parameter configurations in two consecutive stages of Algorithm 3, increasing the node-absorption rates of two bridging nodes from $\delta_{*}$ to $\delta_{**}$ is analogous to removing a community bridge in [38] and increasing the transmission intensities of other nodes (which we call “balancing nodes”) from $\beta_{*}$ to $\beta_{**}$ is analogous to adding edges within a community in [38].
Algorithm 3 SIR simulations on a network $G^{\mathrm{RL}}$ (which we generate using Algorithm 2) for different node-absorption rates (i.e., recovery intensities) and transmission intensities.
@@
Figure 10: The (a) mean outbreak duration, (b) mean final outbreak size, and (c) mean outbreak peak as a function of the stage number of Algorithm 3 with $N_{\rm S}=68$ stages, the graph $G^{\mathrm{RL}}$ that we generate using Algorithm 2 (for $n_{\mathrm{WS}}=12$, $N_{\mathrm{WS}}=20$, and $k_{\mathrm{WS}}=6$), and $N_{\mathrm{sim}}=1000$ simulations of SIR dynamics [for parameter configurations with $\alpha=0.1$, $\delta_{*}=0.2$, $\delta_{**}=1$, $\beta_{*}=0.125$, and the associated value of $\beta_{**}$ that we obtain using (39)]. In (d), we show the map function $L(M_{0})=L(M_{0},P_{e}(D_{\delta}^{(s)},\mathbf{0},t=0.025))$ for the planted partition $M_{0}$ of $G^{\mathrm{RL}}$ as a function the stage number of Algorithm 3 with the same inputs as in (a)–(c), where the node-absorption rates in the diagonal entries of $D_{\delta}^{(s)}$ correspond to the parameter configuration in stage $s$ of Algorithm 3.
@@ 6.2.3 Effective community structure of a single planted community
The 20 ring-lattice subgraphs $G_{1},\ldots,G_{20}$ of the network $G^{\mathrm{RL}}$ in 6.2.1 are isomorphic and randomly connected according to the procedure in Algorithm 2. We now examine the effective community structure of one ring-lattice subgraph of $G^{\mathrm{RL}}$. We use Algorithm 1b with the input $P_{e}(D_{\delta},\mathbf{0},t)$ for the adjacency matrix of $G^{\mathrm{RL}}$, Markov times $t\in(0.01,0.05)$, and node-absorption rates $\delta_{1},\ldots,\delta_{240}$ from the parameter configurations in three stages of Algorithm 3. We use $H=\mathbf{0}$ as an input of Algorithm 1b because the influence of the absorption on the effective community structure is particularly noticeable when $H=\mathbf{0}$ [e.g., see Figures 7(b,d)].
@@
In Figures 11(a) and 11(b), we show the number of communities of $G^{\mathrm{RL}}$ and number of subcommunities of $G_{5}$, respectively, as functions of the Markov time $t$ for partitions that we obtain using Algorithm 1b with the input $P_{e}(D_{\delta},\mathbf{0},t)$, where the node-absorption rates in the diagonal entries of $D_{\delta}$ correspond to the parameter configurations in the initial stage (see the dash-dotted green curve), the stage with the peak duration (see the solid red curve), and the final stage (see the dashed blue curve). In Figures 11(a,b), we observe that the number of communities of the network $G^{\mathrm{RL}}$ [in Figure 11(a)] and the number of subcommunities of the subgraph $G_{5}$ [in Figure 11(b)] increase with the stage number.
@@
Figure 11: Comparison of community structures in the initial, peak-duration, and final stages of Algorithm 3 for SIR dynamics on $G^{\mathrm{RL}}$. (a) The number of communities in $G^{\mathrm{RL}}$ that we obtain using Algorithm 1b with the input $P_{e}(D_{\delta},\mathbf{0},t)$ for the node-absorption rates from the parameter configurations in three stages of Algorithm 3. (b) The number of subcommunities in the planted community $G_{5}$ that we obtain using Algorithm 1b for the node-absorption rates from the parameter configurations in three stages of Algorithm 3. (c,d) The subcommunities of $G_{5}$ that we obtain using Algorithm 1b in (c) the peak-duration stage with $t=0.025$ and (d) the final stage with $t=0.025$. Each color indicates a subcommunity. The node sizes are positively correlated with the node-absorption rates, but they are not linearly proportional to them.
@@ Appendix A Appendix
Let $P_{0}=AW^{-1}$ be the transition-probability matrix of the discrete-time Markov chain that is associated with ~$\tilde{\mathcal{L}}(D_{\delta},\mathbf{0})$, and let $P_{1}=(A+D_{\delta})(W+D_{\delta})^{-1}$ be the transition-probability matrix of the discrete-time Markov chain that is associated with ~$\tilde{\mathcal{L}}(D_{\delta},I)$. Let $\pi$ and $\pi^{\prime}$ be the stationary distributions that are associated with $P_{0}$ and $P_{1}$, respectively. Let $Z_{i}$ be the fundamental matrix that is associated with $P_{i}$ (with $i\in\{1,2\}$). Let $U\mathrel{\mathop{:}}=(1/(\vec{\delta}^{\operatorname{\,T}}\vec{u}))\vec{u}\vec{1}^{\operatorname{\,T}}$ and $\alpha\mathrel{\mathop{:}}=\vec{\delta}^{\operatorname{\,T}}\vec{u}/(\vec{w}^{\operatorname{\,T}}\vec{u}+\vec{\delta}^{\operatorname{\,T}}\vec{u})$, where $\vec{u}=(u_{1},\ldots,u_{n})^{\operatorname{\,T}}$ is a vector in the kernel $\operatorname{Ker}{(W-A)}$ with positive entries $u_{i}$ such that $\sum_{i=1}^{n}u_{i}=1$. We have that
@@
$F_{0}$ $\mathrel{\mathop{:}}=I-P_{0}+\pi\vec{1}^{\operatorname{\,T}}\,,$
$F_{1}$ $\mathrel{\mathop{:}}=I-P_{0}+\pi\vec{1}^{\operatorname{\,T}}-\alpha\pi\vec{1}^{\operatorname{\,T}}\,,$
$F_{2}$ $\mathrel{\mathop{:}}=I-P_{0}+\pi\vec{1}^{\operatorname{\,T}}-\alpha\pi\vec{1}^{\operatorname{\,T}}+\alpha\frac{D_{\delta}\vec{u}}{\vec{\delta}^{\operatorname{\,T}}\vec{u}}\vec{1}^{\operatorname{\,T}}\,,$
$F_{3}$ $\mathrel{\mathop{:}}=I-P_{0}+\pi\vec{1}^{\operatorname{\,T}}-\alpha\pi\vec{1}^{\operatorname{\,T}}+\alpha\frac{D_{\delta}\vec{u}}{\vec{\delta}^{\operatorname{\,T}}\vec{u}}\vec{1}^{\operatorname{\,T}}+\left((1-\alpha)\pi+\alpha\frac{D_{\delta}\vec{u}}{\vec{\delta}^{\operatorname{\,T}}\vec{u}}\right)\vec{1}^{\operatorname{\,T}}D_{\delta}W^{-1}\,.$ (43)
@@
Let ~$\tilde{\mathcal{L}}_{1}\mathrel{\mathop{:}}=\tilde{\mathcal{L}}(D_{\delta},I)=(W-A)(W+D_{\delta})^{-1}$, and let $\vec{d}_{1}$ be the diagonal of $D_{\delta}(W+D_{\delta})^{-1}$. Additionally, let $U\mathrel{\mathop{:}}=\vec{u}\vec{1}^{\operatorname{\,T}}/(\vec{\delta}^{\operatorname{\,T}}\vec{u})$, $U_{1}\mathrel{\mathop{:}}=(W+D_{\delta})U$, and $D_{1}\mathrel{\mathop{:}}=D_{\delta}(W+D_{\delta})^{-1}$. We have that
@@
Let $\vec{d}^{\prime}\mathrel{\mathop{:}}=\vec{d}_{\mathrm{s}}(D_{\delta},I)=\vec{w}+\vec{\delta}=(\omega_{1}+\delta_{1},\ldots,\omega_{n}+\delta_{n})^{\operatorname{\,T}}$ be the scaled rate vector that is associated with the absorption-scaled graph $\tilde{G}(D_{\delta},I)$. With $\mathcal{L}=W-A$, $\alpha\mathrel{\mathop{:}}=\vec{\delta}^{\operatorname{\,T}}\vec{u}/(\vec{w}^{\operatorname{\,T}}\vec{u}+\vec{\delta}^{\operatorname{\,T}}\vec{u})$, $\pi=W\vec{u}/(\vec{w}^{\operatorname{\,T}}\vec{u})$, $Z_{0}=\left(I-AW^{-1}+\pi\vec{1}^{\operatorname{\,T}}\right)^{-1}$, and $Z_{*}=W^{-1}\left(Z_{0}-\pi\vec{1}^{\operatorname{\,T}}\right)$, it follows that
@@
The code that yields the numerical results in Figures 10 and 11 of Section 6 are available at https://gitlab.com/esteban_vargas_bernal/extending-infomap-to-absorbing-random-walks.

### v4

@@ 6.2 Increasing node-absorption rates of bridging nodes instead of removing community bridges
6.2.1 Ring-lattice graphs as planted communities
@@
We use Algorithm 2 to generate the network $G^{\mathrm{RL}}$ (which is inspired by the example in [38] that we described in Section 6.1) on which we simulate an SIR process. [See Figure 10(a) for an example of a subgraph of a network that we generate using Algorithm 2.] We generate a graph $G^{\mathrm{RL}}$ using Algorithm 2 with $N_{\mathrm{WS}}=20$ directed and unweighted ring-lattice subgraphs $G_{1},\ldots,G_{20}$ (the planted communities) of size $n_{\mathrm{WS}}=12$ (i.e., with 12 nodes each) that are connected by community bridges (which we place between node pairs that we choose uniformly at random from all possible pairs with nodes in distinct planted communities). Each node of $G^{\mathrm{RL}}$ has $k_{\mathrm{WS}}=6$ neighbors in its planted community. We choose a small size for the ring-lattice graphs to visualize the effective community structure of one of the 20 ring-lattice subgraphs [see Figures 12(c,d)]. Each of these ring-lattice subgraphs is a one-dimensional lattice with periodic boundary conditions and additional local connections, which we specify in Algorithm 2. In the context of the analogy with the example in [38], each ring-lattice subgraph in our example is a directed WS network in which all edges are bidirectional (i.e., each edge is reciprocated) and the edge-rewiring probability is 00 [27]. We use an edge-rewiring probability of 00 because we seek to examine the effects of changing absorption rates instead of the effects of rewiring edges (which was explored in [38]). The degree of a node of a network $G^{\textrm{RL}}$ that we generate with Algorithm 2 is the random variable $k_{\textrm{WS}}+N_{\textrm{bridges}}$, where $N_{\textrm{bridges}}$ is the number of community bridges that are adjacent to the node. This random variable follows a binomial distribution with parameters $n_{\textrm{WS}}(N_{\textrm{WS}}-1)$ and $2/(n_{\textrm{WS}}(N_{\textrm{WS}}-1))$. The mean degree of $G^{\textrm{RL}}$ is $k_{\textrm{WS}}+2$, and the variance of the degree is $2\left[1-2/(n_{\textrm{WS}}(N_{\textrm{WS}}-1))\right]$. For the graph $G^{\textrm{RL}}$, the standard InfoMap algorithm yields the partition that consists of the planted communities $G_{1},\ldots,G_{N_{\textrm{WS}}}$.
@@
3:  Select $n_{\mathrm{WS}}\times N_{\mathrm{WS}}$ pairs $\{i_{1},i_{2}\}$ of nodes uniformly at random from all node pairs that belong to distinct ring-lattice graphs. Add bidirectional edges between the nodes of each pair. (The pair $\{i_{1},i_{2}\}$ is a “community bridge”, and the nodes $i_{1}$ and $i_{2}$ are “bridging nodes”.)
@@ 6.2.2 Duration, final size, and peak of SIR outbreaks
We use Algorithm 3 to run SIR simulations on the graph $G^{\mathrm{RL}}$ for different node-absorption rates (which correspond to the recovery intensities) and transmission intensities. We refer to a set of node-absorption rates $\delta_{1},\ldots,\delta_{N}$ and transmission intensities $\beta_{1},\ldots,\beta_{N}$ of the $N$ nodes of $G^{\textnormal{RL}}$ as a parameter configuration. In steps 1 and 2 of Algorithm 3, we recursively determine parameter configurations. We refer to each step of this recursive process as a stage. For the parameter configurations in two consecutive stages of Algorithm 3, increasing the node-absorption rates of two bridging nodes from $\delta_{*}$ to $\delta_{**}$ is analogous to removing a community bridge in [38] and increasing the transmission intensities of other nodes (which we call “balancing nodes”) from $\beta_{*}$ to $\beta_{**}$ is analogous to adding edges within a community in [38]. In Figure 10, we show example parameter configurations in two stages of Algorithm 3 for a subgraph of a network that we generate with Algorithm 2.
Algorithm 3 SIR simulations on a network $G^{\mathrm{RL}}$ (which we generate with Algorithm 2) for different node-absorption rates (i.e., recovery intensities) and transmission intensities.
@@
Figure 10: Illustration of a subgraph of a network that we generate with Algorithm 2 and two stages of Algorithm 3 for that network. We show the parameter configurations in stages 1 and 2 of Algorithm 3 for a network $G^{\textrm{RL}}$ from Algorithm 2 with $n_{\textrm{WS}}=6$, $N_{\textrm{WS}}=2$, and $k_{\textrm{WS}}=4$. In this figure, we show only four community bridges (the blue segments) of the $n_{\textrm{WS}}\times N_{\textrm{WS}}=12$ community bridges of $G^{\rm RL}$ between the planted communities $G_{1}=\{1,\ldots,6\}$ and $G_{2}=\{7,\ldots,12\}$. (a) In the parameter configuration in stage 1, all 12 nodes have the same node-absorption rate $\delta_{*}$ and the same transmission intensity $\beta_{*}$. (b) In the parameter configuration in stage 2, the bridging nodes 4 and 7 have the same node-absorption rate $\delta_{**}>\delta_{*}$ and all other nodes have the node-absorption rate $\delta_{*}$. In this stage, nodes 1 and 10 (which are not in the neighborhoods of nodes 4 and 7) have the same transmission intensity $\beta_{**}>\beta_{*}$ and all other nodes have the transmission intensity $\beta_{*}$.
@@
Figure 11: The (a) mean outbreak duration, (b) mean final outbreak size, and (c) mean outbreak peak as a function of the stage number of Algorithm 3 with $N_{\rm S}=68$ stages, the graph $G^{\mathrm{RL}}$ that we generate with Algorithm 2 (for $n_{\mathrm{WS}}=12$, $N_{\mathrm{WS}}=20$, and $k_{\mathrm{WS}}=6$), and $N_{\mathrm{sim}}=1000$ simulations of SIR dynamics [for parameter configurations with $\alpha=0.1$, $\delta_{*}=0.2$, $\delta_{**}=1$, $\beta_{*}=0.125$, and the associated value of $\beta_{**}$ that we obtain using (39)]. In (d), we show the map function $L(M_{0})=L(M_{0},P_{e}(D_{\delta}^{(s)},\mathbf{0},t=0.025))$ for the planted partition $M_{0}$ of $G^{\mathrm{RL}}$ as a function the stage number of Algorithm 3 with the same inputs as in (a)–(c), where the node-absorption rates in the diagonal entries of $D_{\delta}^{(s)}$ correspond to the parameter configuration in stage $s$ of Algorithm 3.
@@ 6.2.3 Effective community structure of a single planted community
The 20 ring-lattice subgraphs $G_{1},\ldots,G_{20}$ of the network $G^{\mathrm{RL}}$ in Section 6.2.1 are isomorphic and randomly connected according to the procedure in Algorithm 2. We now examine the effective community structure of one ring-lattice subgraph of $G^{\mathrm{RL}}$. We use Algorithm 1b with the input $P_{e}(D_{\delta},\mathbf{0},t)$ for the adjacency matrix of $G^{\mathrm{RL}}$, Markov times $t\in(0.01,0.05)$, and node-absorption rates $\delta_{1},\ldots,\delta_{240}$ from the parameter configurations in three stages of Algorithm 3. We use $H=\mathbf{0}$ as an input of Algorithm 1b because the influence of the absorption on the effective community structure is particularly noticeable when $H=\mathbf{0}$ [e.g., see Figures 7(b,d)].
@@
In Figures 12(a) and 12(b), we show the number of communities of $G^{\mathrm{RL}}$ and number of subcommunities of $G_{5}$, respectively, as functions of the Markov time $t$ for partitions that we obtain using Algorithm 1b with the input $P_{e}(D_{\delta},\mathbf{0},t)$, where the node-absorption rates in the diagonal entries of $D_{\delta}$ correspond to the parameter configurations in the initial stage (see the dash-dotted green curve), the stage with the peak duration (see the solid red curve), and the final stage (see the dashed blue curve). In Figures 12(a,b), we observe that the number of communities of the network $G^{\mathrm{RL}}$ [in Figure 12(a)] and the number of subcommunities of the subgraph $G_{5}$ [in Figure 12(b)] increase with the stage number.
@@
Figure 12: Comparison of community structures in the initial, peak-duration, and final stages of Algorithm 3 for SIR dynamics on $G^{\mathrm{RL}}$. (a) The number of communities in $G^{\mathrm{RL}}$ that we obtain using Algorithm 1b with the input $P_{e}(D_{\delta},\mathbf{0},t)$ for the node-absorption rates from the parameter configurations in three stages of Algorithm 3. (b) The number of subcommunities in the planted community $G_{5}$ that we obtain using Algorithm 1b for the node-absorption rates from the parameter configurations in three stages of Algorithm 3. (c,d) The subcommunities of $G_{5}$ that we obtain using Algorithm 1b in (c) the peak-duration stage with $t=0.025$ and (d) the final stage with $t=0.025$. Each color indicates a subcommunity.
@@ Appendix A Appendix
Let $P_{0}=AW^{-1}$ be the transition-probability matrix of the discrete-time Markov chain that is associated with ~$\tilde{\mathcal{L}}(D_{\delta},\mathbf{0})$, and let $P_{1}=(A+D_{\delta})(W+D_{\delta})^{-1}$ be the transition-probability matrix of the discrete-time Markov chain that is associated with ~$\tilde{\mathcal{L}}(D_{\delta},I)$. Let $\pi$ and $\pi^{\prime}$ be the stationary distributions that are associated with $P_{0}$ and $P_{1}$, respectively. Let $Z_{i}$ be the fundamental matrix that is associated with $P_{i}$ (with $i\in\{1,2\}$). Let $U:=(1/(\vec{\delta}^{\operatorname{\,T}}\vec{u}))\vec{u}\vec{1}^{\operatorname{\,T}}$ and $\alpha:=\vec{\delta}^{\operatorname{\,T}}\vec{u}/(\vec{w}^{\operatorname{\,T}}\vec{u}+\vec{\delta}^{\operatorname{\,T}}\vec{u})$, where $\vec{u}=(u_{1},\ldots,u_{n})^{\operatorname{\,T}}$ is a vector in the kernel $\operatorname{Ker}{(W-A)}$ with positive entries $u_{i}$ such that $\sum_{i=1}^{n}u_{i}=1$. We have that
@@
$F_{0}$ $:=I-P_{0}+\pi\vec{1}^{\operatorname{\,T}}\,,$
$F_{1}$ $:=I-P_{0}+\pi\vec{1}^{\operatorname{\,T}}-\alpha\pi\vec{1}^{\operatorname{\,T}}\,,$
$F_{2}$ $:=I-P_{0}+\pi\vec{1}^{\operatorname{\,T}}-\alpha\pi\vec{1}^{\operatorname{\,T}}+\alpha\frac{D_{\delta}\vec{u}}{\vec{\delta}^{\operatorname{\,T}}\vec{u}}\vec{1}^{\operatorname{\,T}}\,,$
$F_{3}$ $:=I-P_{0}+\pi\vec{1}^{\operatorname{\,T}}-\alpha\pi\vec{1}^{\operatorname{\,T}}+\alpha\frac{D_{\delta}\vec{u}}{\vec{\delta}^{\operatorname{\,T}}\vec{u}}\vec{1}^{\operatorname{\,T}}+\left((1-\alpha)\pi+\alpha\frac{D_{\delta}\vec{u}}{\vec{\delta}^{\operatorname{\,T}}\vec{u}}\right)\vec{1}^{\operatorname{\,T}}D_{\delta}W^{-1}\,.$ (43)
@@
Let ~$\tilde{\mathcal{L}}_{1}:=\tilde{\mathcal{L}}(D_{\delta},I)=(W-A)(W+D_{\delta})^{-1}$, and let $\vec{d}_{1}$ be the diagonal of $D_{\delta}(W+D_{\delta})^{-1}$. Additionally, let $U:=\vec{u}\vec{1}^{\operatorname{\,T}}/(\vec{\delta}^{\operatorname{\,T}}\vec{u})$, $U_{1}:=(W+D_{\delta})U$, and $D_{1}:=D_{\delta}(W+D_{\delta})^{-1}$. We have that
@@
Let $\vec{d}^{\prime}:=\vec{d}_{\mathrm{s}}(D_{\delta},I)=\vec{w}+\vec{\delta}=(\omega_{1}+\delta_{1},\ldots,\omega_{n}+\delta_{n})^{\operatorname{\,T}}$ be the scaled rate vector that is associated with the absorption-scaled graph $\tilde{G}(D_{\delta},I)$. With $\mathcal{L}=W-A$, $\alpha:=\vec{\delta}^{\operatorname{\,T}}\vec{u}/(\vec{w}^{\operatorname{\,T}}\vec{u}+\vec{\delta}^{\operatorname{\,T}}\vec{u})$, $\pi=W\vec{u}/(\vec{w}^{\operatorname{\,T}}\vec{u})$, $Z_{0}=\left(I-AW^{-1}+\pi\vec{1}^{\operatorname{\,T}}\right)^{-1}$, and $Z_{*}=W^{-1}\left(Z_{0}-\pi\vec{1}^{\operatorname{\,T}}\right)$, it follows that
@@
The code that yields the numerical results in Figures 11 and 12 of Section 6 are available at https://gitlab.com/esteban_vargas_bernal/extending-infomap-to-absorbing-random-walks.